When 5.00 g of sodium hydrogen carbonate, NaHCO3(s), reacts completely with 70.00 mL of acetic acid, CH3COOH(aq), the temperature increases from 15.5 C to 22.3C. Assume that the acid solution has the same density and specific heat capacity as water and that the mass of the final solution is 70.00 g. Calculate the molar enthalpy of the reaction, Hr of NaHCO3.
1
Expert's answer
2020-05-03T15:09:44-0400

Find the amount of heat released during the reaction:


Q = cm(t2-t1) = 4200*0.07*(22.3 - 15.5) = 1999.2 J


Write the equation for the reaction NaHCO3 and CH3COOH:


NaHCO3(s) + CH3COOH = CH3COONa + H2O + CO2 - "\Delta" Hr


n(NaHCO3) = 5/84 = 0.06 mol. Acetic acid in excess. The enthalpy of the reaction is calculated by the number of moles of NaHCO3.


Q = n(NaHCO3)*Δ Hr, herefrom:


Δ Hr = Q/n(NaHCO3) = 1999.2/0.06 = 33320 J/mol = 33.22 kJ/mol


Molar enthalpy of the reaction is 33.22 kJ/mol.